OpenStudy (anonymous):

What is the solution to the equation? 4(3 + 7) = 4 · x + 4 · 7 A. x = 3 B. x = 4 C. x = 7 D. x = 10

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

No

OpenStudy (skullpatrol):

Use the idea of balance and multiply both sides of the equation by 1/4

OpenStudy (skullpatrol):

$$\Huge 4(3 + 7) = 4 · x + 4 · 7$$ $$\Huge \frac{1}{4}4(3 + 7) = \frac{1}{4}[4 · x + 4 · 7]$$ $$\Huge \frac{1}{\cancel{4}}\cancel{4}(3 + 7) = \frac{1}{\cancel{4}}[\cancel{4} · x + \cancel{4} · 7]$$ $$\Huge 3 + 7= x + 7$$

OpenStudy (anonymous):

Thanks